A. Narrow bridge oscillated up and down 5 times per second.

Period:

Frequency:


B. A tuning fork has a frequency of 252 Hz.

Period:

Frequency:


1
Expert's answer
2021-12-22T09:25:16-0500

(A) Period:


"T=\dfrac{1}{f}=\dfrac{1}{5\ Hz}=0.2\ s."

Frequency:


"f=5\ Hz."

(B) Period:


"T=\dfrac{1}{f}=\dfrac{1}{252\ Hz}=0.004\ s."

Frequency:


"f=252\ Hz."